# Power monoids and their arithmetic: a survey

### v1

@@ Abstract
The non-empty finite subsets of a multiplicatively written monoid form a monoid in their own right, and so do the finite subsets that contain the identity element. Partly due to their unusual arithmetic properties, these structures, known as power monoids, have attracted increasing attention in recent years and have in turn stimulated growing interest in new perspectives in factorization theory, better suited to non-cancellative settings. We survey these developments and briefly review some related aspects.
@@
Let $S$ be a multiplicatively written semigroup (see Sect. 2 for notation and terminology). Endowed with the binary operation of setwise multiplication induced by $S$ on its power set and defined by
@@
Power semigroups provide a natural algebraic framework for a variety of key problems in additive number theory and related areas [6, 47, 46, 19], including Sárközy’s conjecture on the “additive irreducibility” of the set of [non-zero] squares of a finite field of prime order [36, Conjecture 1.6] and Ostmann’s conjecture [29, p. 13] on the “asymptotic additive irreducibility” of the set of (positive rational) primes.
@@
By that time, researchers were mainly interested in the study of varieties (that is, classes of semigroups that are closed under taking homomorphic images, subsemigroups, and finite direct products) generated by $\mathcal{P}(S)$ as $S$ ranges over a specified family of finite semigroups. A catalyst of these early developments was the role of power semigroups in the study of formal languages and automata [3].
@@
Let us just mention here that power semigroups are “strongly non-cancellative structures” (for instance, the large power semigroup of a finite group contains non-trivial idempotents, except in trivial cases), and the classical theory of factorizations is ill-suited to handle such situations. For this reason, in Section 3 we review the basics of an “extended theory of factorizations”, which provides tools to circumvent these and other obstacles inherent in the classical approach.
@@
A turning point in the history of the topic was marked by a paper of Tamura and Shafer [40] that has eventually led to the following questions. Here and throughout, the symbol $\cong$ denotes isomorphism, and every morphism between semigroups is understood to be a semigroup homomorphism.
@@
is itself an isomorphism, called the augmentation of $f$. Moreover, the restriction of $f^{\ast}$ to the non-empty finite subsets of $H$ yields an isomorphism from $\mathcal{P}_{\mathrm{fin}}(H)$ onto $\mathcal{P}_{\mathrm{fin}}(K)$, since clearly $f^{\ast}(X)\in\mathcal{P}_{\mathrm{fin}}(K)$ for all $X\in\mathcal{P}_{\mathrm{fin}}(H)$. See [44, Remark 4] and [19, Section 1] for further details and context.
The answer to Question 1.1(1) is negative for the class of all semigroups [28]; is positive for groups [37], semilattices [27, p. 218], Clifford semigroups [17, Theorem 4.7], cancellative commutative semigroups [44, Corollary 1], etc.; and is open for finite semigroups [24, p. 5], despite some authors having claimed the opposite based on results announced in [39] but never proved.
As for Question 1.1(2), very little seems to be known outside the case where $\mathscr{C}$ is a class of finite semigroups contained in any of the classes that are already addressed by the “positive results” reviewed in the previous paragraph (note that $\mathcal{P}(S)=\mathcal{P}_{\mathrm{fin}}(S)$ if and only if $S$ is finite). More precisely, Bienvenu and Geroldinger have shown in [6, Theorem 3.2(3)] that the question has an affirmative answer for numerical monoids, and the conclusion has been generalized to cancellative commutative semigroups in [44, Corollary 1].
@@
Recent work on power monoids has mainly focused on $\mathcal{P}_{\mathrm{fin},1}(M)$. Most notably, Tringali and Yan [47] posed the following analogue of Questions 1.1, where, as before, the symbol $\cong$ denotes a semigroup isomorphism.
@@
Once again, the core of Question 1.2 lies in proving the “only if ” direction. Indeed, if $f$ is a (semigroup) isomorphism from a monoid $H$ to a monoid $K$, then $f$ maps the identity of $H$ to the identity of $K$ (see, for instance, the last lines of [44, Sec. 2]). Since $f[X]$ is finite for every non-empty finite $X\subseteq H$, it is thus clear that the augmentation of $f$, as defined by Eq. (1), restricts to an isomorphism from $\mathcal{P}_{\mathrm{fin},1}(H)$ to $\mathcal{P}_{\mathrm{fin},1}(K)$.
With these preliminaries in mind, it was shown in [47] that Question 1.2 has a negative answer for left- or right-cancellative monoids (loc. cit., Examples 1.2), and a positive answer for (rational) Puiseux monoids (loc. cit., Theorem 2.5), that is, submonoids of the non-negative rational numbers under addition. The latter result confirmed the “Bienvenu–Geroldinger conjecture”, namely, Conjecture 4.7 in [6].
Rago [32] subsequently proved that Question 1.2 has a negative answer within the class of cancellative commutative monoids. More precisely, he showed that if $H$ and $K$ are commutative valuation monoids with trivial unit groups and isomorphic quotient groups, then $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$. More recently, he characterized all pairs $(H,K)$ of non-isomorphic cancellative commutative monoids such that $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ are isomorphic [30, Theorem 18]. In particular, he obtained in [30, Theorem 14] that if $H$ and $K$ have non-trivial unit groups and $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$, then $H$ is isomorphic to $K$.
@@
The study of automorphisms across a wide range of objects and categories (from graphs to fields, from topological spaces to Banach algebras) is closely connected to many other properties of interest and often provides a deeper understanding of various aspects of their theory.
With these thoughts in mind, we gather from the previous subsection that the augmentation of an automorphism $f$ of a semigroup $S$ restricts to an automorphism of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$. This gives a canonical embedding $\Phi\colon\mathrm{Aut}(S)\hookrightarrow\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$, and it is natural to ask whether $\Phi$ is a (group) isomorphism. Mutatis mutandis, the same question arises for $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(S))$ when $S$ is a monoid.
In the aftermath of their proof [47, Theorem 2.5] of [6, Conjecture 4.7], Tringali and Yan obtained in [46, Theorem 3.2] that the automorphism group of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is cyclic of order two, and conjectured [46, Sec. 4] that $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},0}(H))$ is trivial for every numerical monoid $H\neq\mathbb{N}$. As a follow-up, Tringali and Wen [45, Theorem 3.5] have shown that the automorphism group of the finitary power monoid of $\mathbb{Z}$ (the additive group of integers) is isomorphic to the direct product of a cyclic group of order two and the infinite dihedral group, in stark contrast with the fact that the only non-trivial automorphism of the same group is the map $x\mapsto-x$. In a remarkable feat, Rago [31, Theorem 13] has subsequently proved that, for a finite abelian group $G$, the automorphism group of $\mathcal{P}_{\mathrm{fin},1}(G)$ is (canonically) isomorphic to $\mathrm{Aut}(G)$, except when $G$ is the Klein four-group, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(G))$ is isomorphic to the direct product of two copies of the symmetric group of degree three [31, Example 2]. Even more recently, Wong et al. [49, Theorem 1.3] have established that the automorphism group of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$ of a numerical semigroup $S$ is trivial unless $S=\llbracket k,\infty\rrbracket$ for some integer $k\geq 0$, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$ is cyclic of order two.
@@
To the best of our knowledge, unit-cancellativity was first studied in the commutative setting by Rosales et al. [35, Section 4, p. 72] (under a different name), and in the non-commutative setting by Fan and Tringali [16, Section 2.1]. Acyclicity, on the other hand, was first introduced in [42, Definition 4.2]. Commutative or unit-cancellative monoids are Dedekind-finite [42, Proposition 4.3(i)]. Moreover, every acyclic monoid is unit-cancellative, and the converse holds in the commutative setting.
Given a set $A$, we denote by $\mathscr{F}^{+}(A)$ the free semigroup over $A$ and write $\ast_{A}$ for its operation; when no ambiguity is likely to arise, we drop the subscript $A$ from this notation. As a set, $\mathscr{F}^{+}(A)$ consists of all non-empty finite tuples of elements of $A$, and $\ast_{A}$ is the operation of concatenation of such tuples.
We refer to an element of $\mathscr{F}^{+}(A)$ as an $A$-word, or simply a word when there is no serious risk of confusion. If $\mathfrak{u}$ is an $A$-word, its length is the unique integer $k\geq 1$ such that $\mathfrak{u}$ is a $k$-tuple of elements of $A$. In this case, $\mathfrak{u}=u_{1}\ast\cdots\ast u_{k}$ for some uniquely determined $u_{1},\ldots,u_{k}\in A$, called the letters of $\mathfrak{u}$.
Further notation and terminology, if not explained when first used, are standard or should be clear from the context. In particular, we address the reader to Howie’s monograph [25] for basic aspects of semigroup theory.
@@
Roughly speaking, the theory of factorizations is the study of a broad spectrum of phenomena arising from the possibility or impossibility of extending the Fundamental Theorem of Arithmetic from the integers to more abstract settings where a set comes endowed with a binary operation. The theory has traditionally focused on domains and cancellative monoids, primarily in the commutative setting [21, 20, 22, 4, 38], and the building blocks used in factorizations are atoms, that is, non-units that cannot be written as a product of two other non-units. Only in the past few years have researchers taken the first steps toward a systematic study of factorizations in non-cancellative (and non-commutative) settings, using building blocks that are not necessarily atoms; see [5, Section 2.4] for a review of some earlier scattered results along these lines.
@@
A key feature of these new perspectives is the idea of pairing the monoid $M$ with a preorder (that is, a reflexive and transitive binary relation on $M$) and letting the preorder induce on $M$ a relative notion of irreducibility [42, Section 3]. In this way, the theory gains tremendous flexibility, as neatly explained by Cossu in [9], allowing, for instance, the study of factorizations in non-trivial contexts such as groups, where the classical theory is inapplicable because there are no atoms, or in rings with non-zero zero divisors, where the classical theory loses significance as its invariants tend to blow up in a more or less predictable way.
@@
Atoms (in the sense of the classical theory) and quarks are irreducibles, but the converse need not be true. For instance, the zero element $0_{R}$ of a domain $R$ is irreducible in the multiplicative monoid of $R$, yet it is not an atom, since it is idempotent. Moreover, if $R$ is not a skew field, then $0_{R}$ is not a quark either, because every non-unit $x\in R$ divides $0_{R}$. For a more interesting example, see Theorem 4.7(ii).
@@
Let $a$ be an atom of $M$ (if $M$ has no atoms, the statement is vacuously true), and let $b\in M$ be a divisor of $a$ but not a unit-divisor. By [16, Lemma 2.2(i)], $M$ is Dedekind-finite, and hence $b$ is a non-unit. We need to show that $a\mid_{M}b$.
@@
It should be noted that our distinction between irreducibles and atoms is a delicate point and a common source of confusion. While in the classical theory the two terms are used interchangeably, the former notion is considerably more flexible than the latter and better suited to the extended version of the theory that we consider here; see, e.g., [42, Theorem 3.10], [11, Theorem 5.19], and [10, Theorem 5.1].
@@
A fundamental tenet of factorization theory is that, apart from trivial cases, the arithmetic of a monoid $M$ is controlled by the arithmetic of its proper divisor-closed submonoids. From [10, Corollary 3.6], we know that if $N$ is a divisor-closed submonoid of $M$, then $N$ and $M$ share the same unit-divisors, and every irreducible (resp., atom) of $N$ is also an irreducible (resp., atom) of $M$. We are thus led to the following proposition, which is a (very) special case of [10, Proposition 3.5].
@@
Length sets are among the most studied arithmetic invariants in the classical theory of factorizations. In this regard, it is worth noting that, for an acyclic monoid, every irreducible is an atom [42, Corollary 4.4], so there is no distinction between factorizations and atomic factorizations.
@@
These definitions are modeled on their classical counterparts [21, Theorem 1.1.10.2 and Definitions 1.1.1.4, 1.2.6, 1.3.1, and 1.5.1], and it is fairly obvious that the implications shown in the following diagram hold:
@@
In this section, we survey several results concerning irreducible, atomic, and minimal factorizations in the reduced finitary power monoid $\mathcal{P}_{\mathrm{fin},1}(H)$ of $H$. Related structures, including $\mathcal{P}_{\mathrm{fin},\times}(H)$ and $\mathcal{P}_{\mathrm{fin}}(H)$, will be mentioned only occasionally, since the current state of knowledge does not yet allow for a systematic treatment comparable to that of $\mathcal{P}_{\mathrm{fin},1}(H)$. Some partial results, notably by Fan and Tringali [16], Antoniou and Tringali [5], Bienvenu and Geroldinger [6], Gonzalez et al. [23], and Gotti et al. [1, 14], will not be discussed here.
@@
Thus, $\mathcal{P}_{\mathrm{fin},\times}(H)$ satisfies the ACCP, which completes the proof, since $\mathcal{P}_{\mathrm{fin},\times}(H)$ is Dedekind-finite by (iii) and every Dedekind-finite monoid satisfying the ACCP is known to be factorable [42, Theorem 3.10]. ∎
@@
Obviously, every acyclic monoid is unit-cancellative. Moreover, it is clear that any acyclic monoid $M$ with trivial unit group is torsion-free: if $x^{m}=x^{n}$ for some $x\in M$ and $m,n\in\mathbb{N}^{+}$ with $m<n$, then $x$ must be a unit and hence the identity. It is therefore sufficient to show that $\mathcal{P}_{\mathrm{fin},1}(H)$ is acyclic, as we know from Theorem 4.1(i) that the only unit of $\mathcal{P}_{\mathrm{fin},1}(H)$ is the identity.
@@
For (i), see [42, Proposition 4.11(iii)]. For (ii), see [13, Proposition 2.3], where it is however mistakenly stated (the proof is correct) that $X$ is irreducible but not an atom “if and only if $X=\{1_{H},x\}$ for some $x\in H$ such that $x^{2}=1_{H}$ or $x^{2}=x$” (one should add that $x\neq 1_{H}$). ∎
@@
The equivalence (b) $\Leftrightarrow$ (c) is, on the other hand, a very special case of [10, Theorem 4.11(i)]. It suffices to note that, by Eq. (2), every $X\in\mathcal{P}_{\mathrm{fin},1}(H)$ has only finitely many divisors, and hence $\mathcal{P}_{\mathrm{fin},1}(H)$ is locally of finite type in the sense of [10, Definition 4.1(3)]. For additional details, see [10, Example 4.5(3)]. ∎
@@
We conclude with some results concerning minimal factorizations.
@@
The following conditions hold:
  1. (i)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF-atomic if and only if $H$ is trivial.
  2. (ii)
  3. (iii)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is FmF and BmF.
@@
For (i) and (ii), see Theorem 4.14 and Corollary 4.15 in [5], respectively. As for (i), the conclusion follows directly from [10, Theorem 4.11], in a manner similar to the proof of the equivalence (b) $\Leftrightarrow$ (c) in Theorem 4.15. ∎
For the next result, we recall that a semigroup $S$ is called breakable if $xy=x$ or $xy=y$ for all $x,y\in S$. This is a rather special class of semigroups, first introduced by Rédei [33, Section 27], with a rich structural theory.

### v2

@@ Abstract
The non-empty finite subsets of a multiplicatively written monoid form a monoid in their own right, and so do the finite subsets that contain the identity element. Partly due to their unusual arithmetic properties, these structures, known as power monoids, have attracted increasing attention in recent years and have in turn stimulated growing interest in new perspectives in factorization theory, better suited to non-cancellative or non-commutative settings. We survey these developments and briefly review some related aspects.
@@
Let $S$ be a multiplicatively written semigroup (see Section 2 for notation and terminology). Endowed with the binary operation of setwise multiplication induced by $S$ on its power set and defined by
@@
Power semigroups provide a natural algebraic framework for a variety of key problems in additive number theory and related areas [7, 48, 49, 20], including Sárközy’s conjecture on the “additive irreducibility” of the set of [non-zero] squares of a finite field of prime order [38, Conjecture 1.6] and Ostmann’s conjecture [31, p. 13] on the “asymptotic additive irreducibility” of the set of (positive rational) primes.
@@
By that time, researchers in the area were primarily interested in studying the varieties generated by $\mathcal{P}(S)$ as $S$ ranges over certain families of finite semigroups (here, a variety is a class of semigroups closed under homomorphic images, subsemigroups, and finite direct products). A catalyst of these early developments was the role of power semigroups in the study of formal languages and automata [3].
@@
Let us just mention here that power semigroups are “strongly non-cancellative structures” (for instance, if $G$ is a group, then $XG=GX=G$ for every non-empty $X\subseteq\allowbreak G$), and the classical theory of factorizations is ill-suited to handle such situations. For this reason, in Section 3 we review the basics of an “extended theory of factorizations”, which provides tools to circumvent, or at least mitigate, the side effects arising from these and other obstacles inherent in the classical approach.
@@
A turning point in the history of power semigroups was marked by a paper by Tamura and Shafer [42], eventually leading to the following questions. Here and throughout, the symbol $\cong$ denotes isomorphism, and every morphism between semigroups is understood to be a semigroup homomorphism.
@@
is itself an isomorphism, called the augmentation of $f$. Moreover, the restriction of $f^{\ast}$ to the non-empty finite subsets of $H$ yields an isomorphism from $\mathcal{P}_{\mathrm{fin}}(H)$ onto $\mathcal{P}_{\mathrm{fin}}(K)$, since clearly $f^{\ast}(X)\in\mathcal{P}_{\mathrm{fin}}(K)$ for all $X\in\mathcal{P}_{\mathrm{fin}}(H)$. See [46, Remark 4] and [20, Section 1] for further details and context.
The answer to Question 1.1(1) is negative for the class of all semigroups [30]; is positive for groups [39], semilattices [29, p. 218], Clifford semigroups [18, Theorem 4.7], cancellative commutative semigroups [46, Corollary 1], etc.; and is open for finite semigroups [25, p. 5], despite some authors having claimed otherwise based on unproven results announced in [41]. For a thorough discussion on this point, see
https://mathoverflow.net/questions/508548/
As for Question 1.1(2), very little is known outside the case where $\mathscr{C}$ is a class of finite semigroups contained in any of the classes already covered by the “positive results” reviewed in the previous paragraph. More precisely, Bienvenu and Geroldinger have shown in [7, Theorem 3.2(3)] that the question has an affirmative answer for numerical monoids, and the conclusion has been generalized to cancellative commutative semigroups in [46, Corollary 1].
@@
Recent work on power monoids has mainly focused on $\mathcal{P}_{\mathrm{fin},1}(M)$. Note in this regard that $\mathcal{P}_{\mathrm{fin},1}(M)=\mathcal{P}_{\mathrm{fin},\times}(M)$ if and only if the unit group $M^{\times}$ of $M$ is trivial, while $\mathcal{P}_{\mathrm{fin},\times}(M)=\mathcal{P}(M)$ if and only if $M=M^{\times}$ (that is, $M$ is a group). Consequently, we have that $\mathcal{P}_{\mathrm{fin},1}(M)=\mathcal{P}_{\mathrm{fin}}(M)$ if and only if $M$ is trivial.
Most notably, Tringali and Yan [48] posed the following analogue of Questions 1.1 for $\mathcal{P}_{\mathrm{fin},1}(\cdot)$, where, as before, the symbol $\cong$ denotes a semigroup isomorphism.
@@
Once again, the core of Question 1.2 lies in the “only if ” direction. Indeed, if $f$ is a (semigroup) isomorphism from a monoid $H$ to a monoid $K$, then $f$ maps the identity of $H$ to the identity of $K$ (see, for instance, the last lines of [46, Section 2]). Since $f[X]$ is finite for every non-empty finite $X\subseteq H$, it is thus clear that the augmentation of $f$, as defined by Eq. (1), restricts to an isomorphism from $\mathcal{P}_{\mathrm{fin},1}(H)$ to $\mathcal{P}_{\mathrm{fin},1}(K)$.
With that said, it was shown in [48] that Question 1.2 has a negative answer for left- or right-cancellative monoids (loc. cit., Examples 1.2), and a positive answer for (rational) Puiseux monoids (loc. cit., Theorem 2.5), that is, submonoids of the non-negative rational numbers under addition. The latter result confirmed a conjecture by Bienvenu and Geroldinger [7, Conjecture 4.7].
Rago [34] subsequently proved that Question 1.2 has a negative answer within the class of cancellative commutative monoids. More precisely, he showed that if $H$ and $K$ are commutative valuation monoids with trivial unit groups and isomorphic quotient groups, then $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$. More recently, he characterized all pairs $(H,K)$ of non-isomorphic cancellative commutative monoids such that $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ are isomorphic [32, Theorem 18]. In particular, he obtained that if $H$ and $K$ have non-trivial unit groups and $\mathcal{P}_{\mathrm{fin},1}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin},1}(K)$, then $H$ is isomorphic to $K$ (loc. cit., Theorem 14).
@@
Finally, one could also consider the companion of Question 1.2, obtained by replacing $\mathcal{P}_{\mathrm{fin},1}(\cdot)$ with $\mathcal{P}_{\mathrm{fin},\times}(\cdot)$, leading to the following:
Question 1.3.
Prove or disprove that, for all $H$ and $K$ in a given class of monoids, $\mathcal{P}_{\mathrm{fin},\times}(H)\cong\mathcal{P}_{\mathrm{fin},\times}(K)$ if and only if $H\cong K$.
To date, this has not been studied in the literature, modulo the observation that Questions 1.2 and 1.3 coincide when attention is restricted to classes of monoids with trivial unit group. As a starter, note that, by Theorem 4.2(i), the answer is affirmative for groups, since monoid isomorphisms map units bijectively onto units (recall that any semigroup isomorphism between monoids is a fortiori a monoid isomorphism, i.e., it maps the identity to the identity).
@@
The study of automorphisms for a wide range of objects and categories (from graphs to fields, from topological spaces to operator algebras, and so forth) often yields deeper insight into various aspects of their theory, as it is closely connected with many other properties of interest.
With these thoughts in mind, we gather from the previous subsection that the augmentation of an automorphism $f$ of a semigroup $S$ restricts to an automorphism of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$. This gives a canonical embedding $\Phi\colon\mathrm{Aut}(S)\hookrightarrow\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$, and it is natural to ask whether $\Phi$ is a (group) isomorphism. Mutatis mutandis, the same question arises for $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(M))$ when $M$ is a monoid.
In the aftermath of their proof [48, Theorem 2.5] of [7, Conjecture 4.7], Tringali and Yan obtained in [49, Theorem 3.2] that the automorphism group of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is cyclic of order two, and conjectured [49, Section 4] that $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},0}(H))$ is trivial for every numerical monoid $H\neq\mathbb{N}$. As a follow-up, Tringali and Wen [47, Theorem 3.5] have shown that the automorphism group of the finitary power monoid of $\mathbb{Z}$ (the additive group of integers) is isomorphic to the direct product of a cyclic group of order two and the infinite dihedral group, in stark contrast with the fact that the only non-trivial automorphism of the same group is the map $x\mapsto-x$. In a remarkable feat, Rago [33, Theorem 13] has subsequently proved that, for a finite abelian group $G$, the automorphism group of $\mathcal{P}_{\mathrm{fin},1}(G)$ is (canonically) isomorphic to $\mathrm{Aut}(G)$, except when $G$ is the Klein four-group, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin},1}(G))$ is isomorphic to the direct product of two copies of the symmetric group of degree three [33, Example 2]. Even more recently, Wong et al. [51, Theorem 1.3] have established that the automorphism group of the finitary power semigroup $\mathcal{P}_{\mathrm{fin}}(S)$ of a numerical semigroup $S$ is trivial unless $S=\llbracket k,\infty\rrbracket$ for some integer $k\geq 0$, in which case $\mathrm{Aut}(\mathcal{P}_{\mathrm{fin}}(S))$ is cyclic of order two.
@@
To the best of our knowledge, unit-cancellativity was first studied in the commutative setting by Rosales et al. [37, Section 4, p. 72] (under a different name), and in the non-commutative setting by Fan and Tringali [17, Section 2.1]. Acyclicity, on the other hand, was first introduced in [44, Definition 4.2]. Commutative or unit-cancellative monoids are Dedekind-finite [44, Proposition 4.3(i)]. Moreover, every acyclic monoid is unit-cancellative, and the converse holds in the commutative setting.
Given a set $A$, we denote by $\mathscr{F}^{+}(A)$ the free semigroup over $A$ and write $\ast_{A}$ for its operation; when no ambiguity is likely to arise, we drop the subscript $A$ from this notation. As a set, $\mathscr{F}^{+}(A)$ consists of all non-empty finite tuples of elements of $A$, and $\ast_{A}$ is the operation of concatenation of such tuples. Adding the empty tuple as the identity element for concatenation yields the free monoid $\mathscr{F}(A)$ over $A$.
We refer to an element of $\mathscr{F}(A)$ as an $A$-word, or simply a word when there is no serious risk of confusion. In particular, the empty tuple is called the empty $A$-word. If $\mathfrak{u}$ is a non-empty $A$-word, its length is the unique integer $k\geq 1$ such that $\mathfrak{u}$ is a $k$-tuple of elements of $A$. In this case, $\mathfrak{u}=u_{1}\ast\cdots\ast u_{k}$ for some uniquely determined $u_{1},\ldots,u_{k}\in A$, called the letters of $\mathfrak{u}$. The empty $A$-word has length zero.
Further notation and terminology, if not explained when first used, are standard or should be clear from the context. We address the reader to Howie’s monograph [27] for basic aspects of semigroup theory.
@@
Roughly speaking, the theory of factorizations is the study of a broad spectrum of phenomena arising from the possibility or impossibility of extending the Fundamental Theorem of Arithmetic from the integers to more abstract settings where a set comes endowed with a binary operation. The theory has traditionally focused on domains and cancellative monoids, primarily in the commutative setting [22, 6, 21, 23, 4, 40], and the building blocks used in factorizations are atoms, that is, non-units that cannot be written as a product of two other non-units. Only in the past few years have researchers taken the first steps toward a systematic study of factorizations in non-cancellative (and non-commutative) settings, using building blocks that are not necessarily atoms; see [5, Section 2.4] for a review of some earlier scattered results along these lines.
@@
A key feature of these new perspectives is the idea of pairing the monoid $M$ with a preorder (that is, a reflexive and transitive binary relation on $M$) and letting the preorder induce on $M$ a relative notion of irreducibility [44, Section 3]. In this way, the theory gains tremendous flexibility, as neatly explained by Cossu in [10], allowing, for instance, the study of factorizations in non-trivial contexts such as groups, where the classical theory is inapplicable because there are no atoms, or in rings with non-zero zero divisors, where the classical theory loses significance as its invariants tend to blow up in a more or less predictable way.
@@
Atoms (in the sense of the classical theory) and quarks are irreducibles, but the converse need not be true. For instance, the zero element $0_{R}$ of a domain $R$ is irreducible in the multiplicative monoid of $R$, yet it is not an atom, since it is idempotent. Moreover, if $R$ is not a skew field, then $0_{R}$ is not a quark either, because every non-unit $x\in R$ divides $0_{R}$. For a more interesting example, see Theorem 4.7.
@@
Let $a$ be an atom of $M$ (if $M$ has no atoms, the statement is vacuously true), and let $b\in M$ be a divisor of $a$ but not a unit-divisor. By [17, Lemma 2.2(i)], $M$ is Dedekind-finite, and hence $b$ is a non-unit. We need to show that $a\mid_{M}b$.
@@
It should be noted that our distinction between irreducibles and atoms is a delicate point and a common source of confusion. While in the classical theory the two terms are used interchangeably, the former notion is considerably more flexible than the latter and better suited to the extended version of the theory that we consider here; see, e.g., [44, Theorem 3.10], [12, Theorem 5.19], and [11, Theorem 5.1].
@@
A fundamental tenet of factorization theory is that, apart from trivial cases, the arithmetic of a monoid $M$ is controlled by the arithmetic of its proper divisor-closed submonoids. From [11, Corollary 3.6], we know that if $N$ is a divisor-closed submonoid of $M$, then $N$ and $M$ share the same unit-divisors, and every irreducible (resp., atom) of $N$ is also an irreducible (resp., atom) of $M$. We are thus led to the following proposition, which is a special case of [11, Proposition 3.5].
@@
Length sets are among the most studied arithmetic invariants in the classical theory of factorizations. In this regard, it is worth noting that, for an acyclic monoid, every irreducible is an atom [44, Corollary 4.4], so there is no distinction between factorizations and atomic factorizations.
@@
These definitions are modeled on their classical counterparts [22, Theorem 1.1.10.2 and Definitions 1.1.1.4, 1.2.6, 1.3.1, and 1.5.1], and it is fairly obvious that the implications shown in the following diagram hold:
@@
In this section, we survey several results concerning irreducible, atomic, and minimal factorizations in $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},\times}(H)$, and also prove a number of new ones. The finitary power monoid $\mathcal{P}_{\mathrm{fin}}(H)$ of $H$ will be mentioned only occasionally, since the current state of knowledge does not yet allow for a systematic treatment comparable to that of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},\times}(H)$. Some partial results, notably by Fan and Tringali [17], Antoniou and Tringali [5], Bienvenu and Geroldinger [7], Gonzalez et al. [24], and Gotti et al. [1, 15], will not be discussed here.
@@
Thus, $\mathcal{P}_{\mathrm{fin},\times}(H)$ satisfies the ACCP, which completes the proof, since $\mathcal{P}_{\mathrm{fin},\times}(H)$ is Dedekind-finite by (iii) and every Dedekind-finite monoid satisfying the ACCP is known to be factorable [44, Theorem 3.10]. ∎
@@
Obviously, every acyclic monoid is unit-cancellative. Moreover, it is clear that any acyclic monoid $M$ with trivial unit group is torsion-free: if $x^{m}=x^{n}$ for some $x\in M$ and $m,n\in\mathbb{N}^{+}$ with $m<n$, then $x$ must be a unit and hence the identity. By Theorem 4.1(i), it is therefore sufficient to show that $\mathcal{P}_{\mathrm{fin},1}(H)$ is acyclic.
@@
For (i), see [44, Proposition 4.11(iii)]. For (ii), see [14, Proposition 2.3], where it is however mistakenly stated (the proof is correct) that $X$ is irreducible but not an atom “if and only if $X=\{1_{H},x\}$ for some $x\in H$ such that $x^{2}=1_{H}$ or $x^{2}=x$” (one should add that $x\neq 1_{H}$). ∎
@@
The equivalence (b) $\Leftrightarrow$ (c) is, on the other hand, a special case of [11, Theorem 4.11(i)]. It suffices to note that, by Eq. (2), every $X\in\mathcal{P}_{\mathrm{fin},1}(H)$ has only finitely many divisors (in particular, it is divided by finitely many irreducibles), and hence $\mathcal{P}_{\mathrm{fin},1}(H)$ is locally of finite type in the sense of [11, Definition 4.1(3)]. For additional details, see [11, Example 4.5(3)]. ∎
@@
We finish with some results concerning minimal factorizations, starting with a generalization of [5, Proposition 4.12(i)].
@@
The largest minimal length of a set $X\in\mathcal{P}_{\mathrm{fin},1}(H)$ is smaller than or equal to |X|1|X|-1. In particular, $\mathcal{P}_{\mathrm{fin},1}(H)$ is BmF and FmF.
@@
The equivalence between BmFness and FmFness is a direct consequence of [11, Theorem 4.11(i)], as in the proof of the equivalence between items (b) and (c) in Theorem 4.15. Therefore, we focus below on the first part of the statement.
The claim is trivial if $X=\{1_{H}\}$, when the only factorization of $X$ is the empty word; or if $X$ is irreducible, in which case $|X|$ is at least $2$ and, by Theorem 4.7(ii), $X$ has a unique minimal factorization of length one. So, assume henceforth that $X$ is neither the identity nor an irreducible of $\mathcal{P}_{\mathrm{fin},1}(H)$, and let $\mathfrak{a}=A_{1}\ast\cdots\ast A_{n}$ be a minimal factorization of $X$ (into irreducibles). Note that $n\geq 2$. We claim that
$A_{1}\cdots A_{i}\subsetneq A_{1}\cdots A_{i+1},\quad\text{for every }i\in\llbracket 1,n-1\rrbracket.$ (4)
Indeed, assume the contrary and denote by $\varepsilon$ the empty word. There then exist an index $i\in\llbracket 1,n-1\rrbracket$ such that $A_{1}\cdots A_{i}=A_{1}\cdots A_{i+1}$. Accordingly, set $\mathfrak{b}:=\varepsilon$ if $i=\allowbreak n-1$, and $\mathfrak{b}:=A_{i+2}\ast\cdots\ast A_{n}$ otherwise. Then $A_{1}\ast\cdots\ast A_{i}\ast\mathfrak{b}$ is a proper subword of $\mathfrak{a}$ as well as a factorization of $X$, contradicting the minimality of $\mathfrak{a}$.
Consequently, Eq. (4) holds, and hence $2\leq|A_{1}\cdots A_{i}|<|A_{1}\cdots A_{i+1}|\leq|X|$ for all $i\in\llbracket 1,n-1\rrbracket$, which implies at once that $n\leq|X|-1$. ∎
For the next theorem, we recall that a semigroup $S$ is breakable if $xy\in\{x,y\}$ for all $x,y\in S$. This is a rather special class of semigroups, first introduced by Rédei [35, Section 27], with a rich structural theory.